# A practical identifiability criterion leveraging weak-form parameter estimation

### v3

@@ Abstract
In this work, we define a practical identifiability criterion, $(e,q)$-identifiability, based on a parameter $e$, reflecting the noise in observed variables, and a parameter $q$, reflecting the mean-square error of the parameter estimator. This criterion is better able to encompass changes in the quality of the parameter estimate(s) due to increased noise in the data (compared to existing criteria based solely on average relative errors). We illustrate the usefulness of the criteria in several challenging identifiability studies, involving parameter estimation in partially observed systems. Furthermore, we leverage a weak-form equation error-based method of parameter estimation for systems with unobserved variables to assess practical identifiability far more quickly in comparison to output error-based parameter estimation. We do so by generating weak-form input-output equations using differential algebra techniques, as previously proposed by Boulier et al [1], and then applying Weak form Estimation of Nonlinear Dynamics (WENDy) to obtain parameter estimates. This method is computationally efficient and robust to noise, as demonstrated through two classical biological modeling examples.
@@ 1 Introduction
Parameter estimation is central to mathematical modeling in the biological sciences. However, despite the increasing sophistication of computational and statistical techniques over the years, there are many examples of researchers using the same model and obtaining widely varying parameter estimates for the same biological phenomena [2, 3, 4, 5, 6]. It is well-recognized that parameter estimation in dynamical systems remains a challenging task [7]. While some of this variation is due to causes such as quality and quantity of data and uncertainty in initial conditions, the primary reason for that difficulty lies in the relationship between the model structure and data. Analyzing model identifiability, i.e., the ability to recover unique parameters of a system from observations, is thus critical to developing robust and replicable inference in biological systems.
@@
Structural identifiability addresses whether model parameters can be uniquely determined given the observation of some output variables of a system. There are multiple approaches to assess the structural identifiability of a model including but not limited to a Taylor series expansion [8, 9, 10], the implicit function theorem [9, 10], generalizations of observability criteria [11, 12, 13], and differential algebra [^1] A subfield of algebra created by Ritt and his student Kolchin to study the existence of solutions to differential equations ([14, 15]). approaches [16, 17, 18, 19].
@@
However, in practice, the quantity, frequency, and timing of available data have been shown to have large impacts on the reliability of parameter estimates. For examples in models of acute viral infections, we direct the reader to [2, 20, 21]. Furthermore, choices in the numerical discretizations (needed to approximate solutions) can also impact parameter estimates [22, 23]. Practical identifiability [^2] This has also been referred to as statistical identifiability. refers to whether parameters can be uniquely estimated given the available observations and the chosen parameter estimation method, including numerics, and is directly related to uncertainty quantification of model parameters. Practical identifiability methods can be applied to study estimator properties for a model with a specific observed dataset (a posteriori), or to study general properties of a model in different hypothetical data scenarios, i.e., noise and collection frequency (a priori). The latter can provide important insights to guide experimental design, power analyses, and data collection.
@@
While simple models can benefit from traditional statistical methods, such as the Fisher Information Matrix, most practical identifiability techniques for complex models rely on repeated simulations, which can become computationally expensive. In contrast, weak-form based parameter estimation, which is typically much faster than output error-based methods, can address the challenge incurred by repeating the parameter estimation. The idea of using the weak form in system identification can be traced back to the Equations-of-Motion method described in [24], in which an equation (in the strong form) is multiplied by a compactly supported test function [^3] Shinbrot calls them method functions $\phi$ and then the equation is integrated. As the derivatives of the test functions are also assumed to have compact support, integration by parts removes the need to evaluate the derivatives of the data. Shinbrot proposed that the $\phi$’s be trigonometric functions with different frequencies, while Loeb and Cahen [25] later proposed piecewise polynomials (denoting their approach as the Modulating Functions Method). Recently, our group has developed a suite of weak form-based parameter estimation and equation learning methods (WENDy and WSINDy) in the spirit of Shinbrot, Loeb and Cahen, but with mathematically motivated test functions [26, 27, 28]. This choice of test functions improves both the computational efficiency of performing parameter estimation and is highly robust to noise in the data. These weak form-based methods have also been applied to biological and environmental systems [29, 30, 31].
@@
One challenge in using WENDy type weak-form estimation methods is that, thus far, all state variables had to be measured. However, in biological systems, data is rarely observed for all model compartments. For example, epidemiological models are often calibrated to only a subset of model compartments, such as mortality or hospitalization data. Accordingly, in this study, we consider differential elimination methods to collapse a system of low-order equations to an equivalent lower-dimensional high-order system with only observable variables. This technique has been applied in [32, 33, 1, 34], and, in some cases, these systems have been cast into either integral or weak forms [1, 34]. However, previous applications of these elimination methods have not yielded a robust weak-form parameter estimation process.
@@
In this work, we propose a process for efficiently assessing a priori model identifiability by leveraging weak-form parameter estimation. First, we apply differential elimination methods to generate an input-output equation, for the states which will be observed. At this step, we can also assess the structural identifiability of the model, which is a prerequisite to practical identifiability. Second, we cast the input-output equation into its weak form. Third, we perform weak-form parameter estimation on repeated simulations of noisy data. Note, unlike output-error based methods, this method is feasible even for large sample sizes and high levels of measurement error when the Weak form Estimation of Nonlinear Dynamics (WENDy) is used for parameter estimation, due to the method’s robustness to noise and computational efficiency. Finally, we assess practical identifiability from these samples using a newly defined criterion, which we refer to as $(e,q)$-identifiability.
@@ 2.1 Structural Identifiability
One way of assessing the structural identifiability of a system of differential equations is by studying the input-output equation. Input-output equations have been used to study the structural identifiability of numerous biological systems, for examples, see [35, 36, 21, 37, 38]. An input-output equation is a differential equation that depends only on the inputs, outputs, their respective derivatives, and the model parameters. Given an input-output equation for a system of differential equations, global structural identifiability can be restated as given below.
@@
A model is said to be globally structurally identifiable if coefficient map of the corresponding input-output equation, $c(w)$ is injective. [^4] Local structural identifiability can be similarly restated. [^5] For this definition to be equivalent to Definition 1, there is an criterion on the Wronskian of the differential monomials from the input-output equation. For more details, see [39].
Input-output equations can be found from a system of differential equations using differential elimination algorithms, including the Ritt, the Ritt-Kolchin [15], and the Rosenfeld-Groebner [40] algorithms, or with a more recently developed, projection-based algorithm [19]. There are many computational tools available to obtain input-output equations, including [18, 41, 42]. In this study, we use the DifferentialAlgebra v4 package [43] in Python to generate input-output equations, which applies a Rosenfeld-Groebner algorithm. Note, an input-output equation may not exist or may not be computationally feasible to discover; however, addressing the weak-form estimation in such cases is a subject of future work. This paper focuses on a weak-form method for the case that an input-output equation, as referred to in Definition 5, can be found.
@@ 2.2 Practical Identifiability
There are many proposed methods to assess the practical identifiability of a system, and there is no commonly agreed-upon single criterion to determine practical identifiability. The Fisher Information Matrix (FIM) can be used to approximate the parameter covariance matrix and as a measure of local practical identifiability [44, 45], however, this criterion is difficult to calculate in more complex systems where likelihood modes may be difficult to find. The FIM method has been used to assess the practical identifiability in biological systems [35, 3]. In these more complicated systems, Profile Likelihood based methods test practical identifiability through successive refittings in the parameter space, which give insight into the shape of the likelihood and whether it is uniquely maximized [^6] Depending on the choice of estimator or proxy used for the likelihood, this problem may be restated as a minimization. [46]. Profile likelihood methods have been used to assess practical identifiability in biological systems as well [47, 37].
Simulation-based methods of assessing practical identifiability involve repeated estimation performed using a prescribed error model for the observed data. From these repeated estimations, one criterion that has been proposed to assess practical identifiability is the average relative error. Examples of applications of this criterion applied to biological systems include [48, 21]. Simulation-based methods can be applied to any choice of objective function, but they also require numerous repetitions of the chosen parameter estimation method, which can become challenging to implement if the estimation method is not computationally efficient.
@@
A parameter, $w_{i}$, is said to be $(e,q)$-identifiable if, for a given observation error ratio, $e$, the mean-squared error (MSE) of the parameter estimator, $\widehat{w}_{i}$, is below the maximum tolerated mean-squared error, $M_{i}$, i.e., $MSE(\widehat{w}_{i}(\mathbf{e}))<(\mathbf{q}{\color[rgb]{0,0,0}w_{i}})^{2}~.$ Furthermore, a model is said to be $(e,q)$-identifiable if the above holds for all $i$.
@@ 3.2 Weak form Estimation of Nonlinear Dynamics (WENDy)
Recently, our group has developed the Weak form Estimation of Nonlinear Dynamics (WENDy) method for parameter estimation [28]. Briefly, in the WENDy method, a system is cast in its weak form, then data are substituted directly into the equation, and parameters are estimated by solving a regression problem. A more detailed description of the method can be found in [28]. Most importantly, this method allows us to estimate parameters from systems in the form of (2) directly using WENDy [^7] Note that in this formulation, we are limited to problems that are linear-in-parameters (not to be confused with linear problems). Recently, [49] has addressed problems that are nonlinear-in-parameters..
@@
Here $\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\Phi_{k}:=[\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\phi_{k}(t_{0})|...|\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\phi_{k}(t_{M})]\mathcal{Q}$ where $\mathcal{Q}$ is a diagonal matrix containing the Newton-Cotes quadrature weights. Let $\mathbf{C_{\phi,y}}=\sum_{i=0}^{n}(-1)^{i}(\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\Phi)C_{i}(\mathbf{y})$ and $\mathbf{H_{\phi,y}}=\sum_{i=0}^{n}(-1)^{i+1}(\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\Phi)H_{i}(\mathbf{y})$. By minimizing the contribution of integration error to the residual, this discretization leads to the following corresponding linear system $\mathbf{G}\mathbf{w}^{\star}\approx\mathbf{b}$ [28], where
@@
As presented in [28], the WENDy algorithm addresses the Errors-In-Variable nature of the linearized problem using iterative reweighting to approximate the correct underlying covariance structure. This reweighting is dependent on the assumed noise structure of the observed data. For Gaussian additive noise, we apply the covariance correction using Algorithm 2 in [28]. Additionally, it has been shown in [49] that, through a strategic choice of transformations to the observations $\mathbf{y}$, it is possible to approximate the covariance structure for systems where observational noise is multiplicative and lognormal (i.e. $\mathbf{y}=\mathbf{y}^{\star}\varepsilon$ where $\log(\varepsilon(\mathbf{t}))\sim\mathcal{N}(0,\sigma^{2})$).

### v4

@@ Abstract
In this work, we define a practical identifiability criterion, $(e,q)$-identifiability, based on a parameter $e$, reflecting the noise in observed variables, and a parameter $q$, reflecting the mean-square error of the parameter estimator. This criterion is better able to encompass changes in the quality of the parameter estimate(s) due to increased noise in the data (compared to existing criteria based solely on average relative errors). We illustrate the usefulness of the criteria in several challenging identifiability studies, involving parameter estimation in partially observed systems. Furthermore, we leverage a weak-form equation error-based method of parameter estimation for systems with unobserved variables to assess practical identifiability far more quickly in comparison to output error-based parameter estimation. We do so by generating weak-form input-output equations using differential algebra techniques, as previously proposed by Boulier et al [BoulierKorporalLemaireEtAl2014ComputerAlgebrainScientificComputing], and then applying Weak form Estimation of Nonlinear Dynamics (WENDy) to obtain parameter estimates. This method is computationally efficient and robust to noise, as demonstrated through two classical biological modeling examples.
@@ 1 Introduction
Parameter estimation is central to mathematical modeling in the biological sciences. However, despite the increasing sophistication of computational and statistical techniques over the years, there are many examples of researchers using the same model and obtaining widely varying parameter estimates for the same biological phenomena [NguyenKlawonnMikolajczykEtAl2016PLoSONE, GoncalvesMentreLemenuel-DiotEtAl2020AAPSJ, StepaniantsHastewellSkinnerEtAl2023arXiv230404818, SimpsonMaclaren2024BullMathBiol, ZitzmannKeRibeiroEtAl2024PLoSComputBiola]. It is well-recognized that parameter estimation in dynamical systems remains a challenging task [McGoffMukherjeePillai2015StatistSurva]. While some of this variation is due to causes such as quality and quantity of data and uncertainty in initial conditions, the primary reason for that difficulty lies in the relationship between the model structure and data. Analyzing model identifiability, i.e., the ability to recover unique parameters of a system from observations, is thus critical to developing robust and replicable inference in biological systems.
@@
Structural identifiability addresses whether model parameters can be uniquely determined given the observation of some output variables of a system. There are multiple approaches to assess the structural identifiability of a model including but not limited to a Taylor series expansion [WielandHauberRosenblattEtAl2021CurrentOpinioninSystemsBiology, ChisBangaBalsa-Canto2011IFACProceedingsVolumes, MiaoWuXue2014JournaloftheAmericanStatisticalAssociation], the implicit function theorem [ChisBangaBalsa-Canto2011IFACProceedingsVolumes, MiaoWuXue2014JournaloftheAmericanStatisticalAssociation], generalizations of observability criteria [StigterMolenaar2015Automatica, VanWilligenburgStigterMolenaar2022NonlinearDyn, VillaverdeBarreiroPapachristodoulou2016PLoSComputBiol], and differential algebra [^1] A subfield of algebra created by Ritt and his student Kolchin to study the existence of solutions to differential equations ([Ritt1950, Kolchin1985]). approaches [LjungGlad1994Automatica, MeshkatEisenbergDiStefano2009MathematicalBiosciences, BelluSaccomaniAudolyEtAl2007ComputerMethodsandProgramsinBiomedicine, DongGoodbrakeHarringtonEtAl2023SIAMJApplAlgebraGeometry].
@@
However, in practice, the quantity, frequency, and timing of available data have been shown to have large impacts on the reliability of parameter estimates. For examples in models of acute viral infections, we direct the reader to [NguyenKlawonnMikolajczykEtAl2016PLoSONE, CiupeTuncer2022SciRep, Heitzman-BreenLiyanageDuggalEtAl2024RSocOpenSci]. Furthermore, choices in the numerical discretizations (needed to approximate solutions) can also impact parameter estimates [DukicBortz2018InverseProblSciEng, NardiniBortz2019InverseProbl]. Practical identifiability [^2] This has also been referred to as statistical identifiability. refers to whether parameters can be uniquely estimated given the available observations and the chosen parameter estimation method, including numerics, and is directly related to uncertainty quantification of model parameters. Practical identifiability methods can be applied to study estimator properties for a model with a specific observed dataset (a posteriori), or to study general properties of a model in different hypothetical data scenarios, i.e., noise and collection frequency (a priori). The latter can provide important insights to guide experimental design, power analyses, and data collection.
@@
While simple models can benefit from traditional statistical methods, such as the Fisher Information Matrix, most practical identifiability techniques for complex models rely on repeated simulations, which can become computationally expensive. In contrast, weak-form based parameter estimation, which is typically much faster than output error-based methods, can address the challenge incurred by repeating the parameter estimation. The idea of using the weak form in system identification can be traced back to the Equations-of-Motion method described in [Shinbrot1954NACATN3288], in which an equation (in the strong form) is multiplied by a compactly supported test function [^3] Shinbrot calls them method functions $\phi$ and then the equation is integrated. As the derivatives of the test functions are also assumed to have compact support, integration by parts removes the need to evaluate the derivatives of the data. Shinbrot proposed that the $\phi$’s be trigonometric functions with different frequencies, while Loeb and Cahen [LoebCahen1963Automatisme] later proposed piecewise polynomials (denoting their approach as the Modulating Functions Method). Recently, our group has developed a suite of weak form-based parameter estimation and equation learning methods (WENDy and WSINDy) in the spirit of Shinbrot, Loeb and Cahen, but with mathematically motivated test functions [MessengerBortz2021JComputPhys, MessengerBortz2021MultiscaleModelSimul, BortzMessengerDukic2023BullMathBiol]. This choice of test functions improves both the computational efficiency of performing parameter estimation and is highly robust to noise in the data. These weak form-based methods have also been applied to biological and environmental systems [MessengerWheelerLiuEtAl2022JRSocInterface, MessengerDwyerDukic2024JRSocInterface, MinorMessengerDukicEtAl2025JGeophysResMachLearnComput]. Morevoer, we have also performed anlayses establishing asymptotic properties of WSINDy [MessengerBortz2025IMAJNumerAnal] as well as bias and coverage properties of WENDy [ChawlaBortzDukic2025HandbookofVisualExperimentalandComputationalMathematicsBridgesthroughData]. For an overview of this novel class of methods, we direct the interested reader to [MessengerTranDukicEtAl2024SIAMNews, BortzMessengerTran2024NumericalAnalysisMeetsMachineLearning].
@@
One challenge in using WENDy type weak-form estimation methods is that, thus far, all state variables had to be measured. However, in biological systems, data is rarely observed for all model compartments. For example, epidemiological models are often calibrated to only a subset of model compartments, such as mortality or hospitalization data. Accordingly, in this study, we consider differential elimination methods to collapse a system of low-order equations to an equivalent lower-dimensional high-order system with only observable variables. This technique has been applied in [Denis-VidalJoly-BlanchardNoiret2003NumericalAlgorithms, VerdiereDenis-VidalJoly-BlanchardEtAl2005IntJApplMathComputSci, BoulierKorporalLemaireEtAl2014ComputerAlgebrainScientificComputing, VerdiereZhuDenis-Vidal2018JournalofComputationalandAppliedMathematics], and, in some cases, these systems have been cast into either integral or weak forms [BoulierKorporalLemaireEtAl2014ComputerAlgebrainScientificComputing, VerdiereZhuDenis-Vidal2018JournalofComputationalandAppliedMathematics, WongvanichHannSirisena2015MathematicalBiosciences]. However, previous applications of these elimination methods have not yielded a robust weak-form parameter estimation process.
@@
In this work, we propose a process for efficiently assessing a priori model identifiability by leveraging weak-form parameter estimation. First, we apply differential elimination methods to generate an input-output equation, for the states which will be observed. At this step, we can also assess the structural identifiability of the model, which is a prerequisite to practical identifiability. Second, we cast the input-output equation into its weak form. Third, we perform weak-form parameter estimation on repeated simulations of noisy data. Note, unlike output error based methods, this method is feasible even for large sample sizes and high levels of measurement error when the Weak form Estimation of Nonlinear Dynamics (WENDy) is used for parameter estimation, due to the method’s robustness to noise and computational efficiency. Finally, we assess practical identifiability from these simulations using a newly defined criterion, which we refer to as $(e,q)$-identifiability.
@@ 2.1 Structural Identifiability
One way of assessing the structural identifiability of a system of differential equations is by studying the input-output equation. Input-output equations have been used to study the structural identifiability of numerous biological systems, for examples, see [EisenbergRobertsonTien2013JournalofTheoreticalBiology, ChowellDahalLiyanageEtAl2023JMathBiol, Heitzman-BreenLiyanageDuggalEtAl2024RSocOpenSci, LiyanageHeitzman-BreenTuncerEtAl2024MBEa, LiyanageChowellPogudinEtAl2025Viruses]. An input-output equation is a differential equation that depends only on the inputs, outputs, their respective derivatives, and the model parameters. Given an input-output equation for a system of differential equations, global structural identifiability can be restated as given below.
@@
A model is said to be globally structurally identifiable if coefficient map of the corresponding input-output equation, $c(w)$ is injective. [^4] Local structural identifiability can be similarly restated. [^5] For this definition to be equivalent to Definition 1, there is an criterion on the Wronskian of the differential monomials from the input-output equation. For more details, see [OvchinnikovPogudinThompson2023AAECC].
Input-output equations can be found from a system of differential equations using differential elimination algorithms, including the Ritt, the Ritt-Kolchin [Kolchin1985], and the Rosenfeld-Groebner [Boulier2007GrobnerBasesinSymbolicAnalysis] algorithms, or with a more recently developed, projection-based algorithm [DongGoodbrakeHarringtonEtAl2023SIAMJApplAlgebraGeometry]. There are many computational tools available to obtain input-output equations, including [BelluSaccomaniAudolyEtAl2007ComputerMethodsandProgramsinBiomedicine, MeshkatKuoDiStefano2014PLoSONE, HongOvchinnikovPogudinEtAl2019Bioinformatics]. In this study, we use the DifferentialAlgebra v4 package [BoulierThiery2024] in Python to generate input-output equations, which applies a Rosenfeld-Groebner algorithm. Note, an input-output equation may not exist or may not be computationally feasible to discover; however, addressing the weak-form estimation in such cases is a subject of future work. This paper focuses on a weak-form method for the case that an input-output equation, as referred to in Definition 5, can be found.
@@ 2.2 Practical Identifiability
There are many proposed methods to assess the practical identifiability of a system, and there is no commonly agreed-upon single criterion to determine practical identifiability. The Fisher Information Matrix (FIM) can be used to approximate the parameter covariance matrix and as a measure of local practical identifiability [JacquezGreif1985MathematicalBiosciences, LandawDiStefano1984AmericanJournalofPhysiology-RegulatoryIntegrativeandComparativePhysiology], however, this criterion is difficult to calculate in more complex systems where likelihood modes may be difficult to find. The FIM method has been used to assess the practical identifiability in biological systems [EisenbergRobertsonTien2013JournalofTheoreticalBiology, GoncalvesMentreLemenuel-DiotEtAl2020AAPSJ]. In these more complicated systems, Profile Likelihood based methods test practical identifiability through successive refittings in the parameter space, which give insight into the shape of the likelihood and whether it is uniquely maximized [^6] Depending on the choice of estimator or proxy used for the likelihood, this problem may be restated as a minimization. [MurphyVanDerVaart2000JournaloftheAmericanStatisticalAssociation]. Profile likelihood methods have been used to assess practical identifiability in biological systems as well [KreutzRaueKaschekEtAl2013TheFEBSJournal, LiyanageHeitzman-BreenTuncerEtAl2024MBEa].
Simulation-based methods of assessing practical identifiability involve repeated estimation performed using a prescribed error model for the observed data. From these repeated estimations, one criterion that has been proposed to assess practical identifiability is the average relative error. Examples of applications of this criterion applied to biological systems include [TuncerMarcthevaLaBarreEtAl2018BullMathBiol, Heitzman-BreenLiyanageDuggalEtAl2024RSocOpenSci]. Simulation-based methods can be applied to any choice of objective function, but they also require numerous repetitions of the chosen parameter estimation method, which can become challenging to implement if the estimation method is not computationally efficient.
@@ 2.3 $(e,q)$-identifiability
A parameter, $w_{i}$, is said to be $(e,q)$-identifiable if, for a given observation error ratio, $e$, the mean-squared error (MSE) of the parameter estimator, $\widehat{w}_{i}$, is below the maximum tolerated mean-squared error, $M_{i}$, i.e., $MSE(\widehat{w}_{i}({\color[rgb]{0,0,0}e}))<({\color[rgb]{0,0,0}q}{\color[rgb]{0,0,0}w_{i}})^{2}~.$ Furthermore, a model is said to be $(e,q)$-identifiable if the above holds for all $i$.
@@ 3.2 Weak form Estimation of Nonlinear Dynamics (WENDy)
Recently, our group has developed the Weak form Estimation of Nonlinear Dynamics (WENDy) method for parameter estimation [BortzMessengerDukic2023BullMathBiol]. Briefly, in the WENDy method, a system is cast in its weak form, then data are substituted directly into the equation, and parameters are estimated by solving a regression problem. A more detailed description of the method can be found in [BortzMessengerDukic2023BullMathBiol]. Most importantly, this method allows us to estimate parameters from systems in the form of (2) directly using WENDy [^7] Note that in this formulation, we are limited to problems that are linear-in-parameters (not to be confused with linear problems). Recently, [RummelMessengerBeckerEtAl2025arXiv250208881] has addressed problems that are nonlinear-in-parameters..
@@
Here $\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\Phi_{k}:=[\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\phi_{k}(t_{0})|...|\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\phi_{k}(t_{M})]\mathcal{Q}$ where $\mathcal{Q}$ is a diagonal matrix containing the Newton-Cotes quadrature weights. Let $\mathbf{C_{\phi,y}}=\sum_{i=0}^{n}(-1)^{i}(\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\Phi)C_{i}(\mathbf{y})$ and $\mathbf{H_{\phi,y}}=\sum_{i=0}^{n}(-1)^{i+1}(\frac{\textsf{d}^{i}}{\textsf{d}t^{i}}\Phi)H_{i}(\mathbf{y})$. By minimizing the contribution of integration error to the residual, this discretization leads to the following corresponding linear system $\mathbf{G}\mathbf{w}^{\star}\approx\mathbf{b}$ [BortzMessengerDukic2023BullMathBiol], where
@@
As presented in [BortzMessengerDukic2023BullMathBiol], the WENDy algorithm addresses the Errors-In-Variable nature of the linearized problem using iterative reweighting to approximate the correct underlying covariance structure [^8] The approximated covariance can also be used to calculate parameter confidence intervals. For a derivation of the parameter covariance approximation, we direct readers to [BortzMessengerDukic2023BullMathBiol].. This reweighting is dependent on the assumed noise structure of the observed data. For Gaussian additive noise, we apply the covariance correction using Algorithm 2 in [BortzMessengerDukic2023BullMathBiol]. Additionally, it has been shown in [RummelMessengerBeckerEtAl2025arXiv250208881] that, through a strategic choice of transformations to the observations $\mathbf{y}$, it is possible to approximate the covariance structure for systems where observational noise is multiplicative and lognormal (i.e. $\mathbf{y}=\mathbf{y}^{\star}\varepsilon$ where $\log(\varepsilon(\mathbf{t}))\sim\mathcal{N}(0,\sigma^{2})$).